# Heterophily-Aware Fair Recommendation using Graph Convolutional Networks

### v2

@@ Abstract
In recent years, graph neural networks (GNNs) have become a popular tool to improve the accuracy and performance of recommender systems. Modern recommender systems are not only designed to serve end users, but also to benefit other participants, such as items and items providers. These participants may have different or conflicting goals and interests, which raise the need for fairness and popularity bias considerations. GNN-based recommendation methods also face the challenges of unfairness and popularity bias and their normalization and aggregation processes suffer from these challenges. In this paper, we propose a fair GNN-based recommender system, called HetroFair, to improve items’ side fairness. HetroFair uses two separate components to generate fairness-aware embeddings: i) fairness-aware attention which incorporates dot product in the normalization process of GNNs, to decrease the effect of nodes’ degrees, and ii) heterophily feature weighting to assign distinct weights to different features during the aggregation process. In order to evaluate the effectiveness of HetroFair, we conduct extensive experiments over six real-world datasets. Our experimental results reveal that HetroFair not only alleviates the unfairness and popularity bias on items’ side, but also achieves superior accuracy on users’ side. Our implementation is publicly available at https://github.com/NematGH/HetroFair.
@@ 1 Introduction
Graph-based recommendation has become more popular in recent years, due to the existence of a bipartite graph in the nature of recommender systems [5, 14, 25, 29, 30, 32]. In this bipartite graph, nodes are divided into two sets of users and items, and edges are ratings or interactions between them. Graph neural networks (GNNs) are a powerful method for deep learning in graph-structured data, as they can capture the complex and non-Euclidean relationships between the nodes and edges of a graph [43, 10, 6, 45, 19]. With the widespread use of graph neural networks in recommender systems, the competition to build systems with higher accuracy entered a new phase [14, 9, 18, 37]. Early methods attempted to improve the accuracy of recommender systems, to better satisfy end users. However, users are not the only stakeholders and accuracy is not the sole objective. Recent researches demonstrate that recommender systems are multi-stakeholder systems in which in addition to end users, other participants (especially items and items providers) should also be considered, during the recommendation process [1, 2, 4, 21, 23]. On the other hand, the power law characteristic of the recommendation graphs shows that a large number of items possess a low degree, while only a few of them have a large degree. Figure 1 illustrates the degree distribution of two real-world recommendation datasets. This feature causes a phenomenon, known in the literature as popularity bias, where popular items are recommended more frequent than the other items. As a result, items with few interactions (long-tail items) are ignored, even if they are of high quality and relevance [3, 20]. Similar investigations have been done for GNN-based recommendation, revealing that long-tail items are often sacrificed during message normalization and aggregation of GNNs [39, 42].
Various methods have been proposed in recent years for long-tail recommendation, which aim to recommend items that belong to the long tail of the distribution of ratings or popularity and at the same time, try to achieve high accuracy on the users’ side. The authors of [44] introduce regularization techniques to overcome the popularity bias. Some other works concentrate on creating multiple graphs for different objectives and simultaneously optimize those graphs in an integrated manner [17, 28]. A few other works directly target the normalization and aggregation processes in the recommendation systems that utilize graph neural networks [39, 42]. However, these GNN-based recommenders either consider static weights for message normalization, or ignore different aspects of items/users.
@@
In this paper, we propose the HetroFair model, to improve items’ side fairness in GNN-based recommendation systems. To do so, first we directly address the cause of popularity bias and reduce the effect of nodes’ degrees on enlarging the values of their embeddings, via involving dot products during the normalization process. Second, with respect to the heterophily property in graphs, we distinguish between different aspects of an item, by considering an specific weight for each feature during the aggregation process. In the heterophily property, nodes of the graph that are not similar to each other, tend to be connected to each other. To the best of our knowledge, our model is the first approach that addresses fairness in GNN-based recommendation using the heterophily property. Our key contributions are summarized as follows:
@@
  • We theoretically analyze the effect of message normalization in GNN-based recommendation on popularity bias.
@@
  • To generate fair embeddings for users and items, we design a fairness-aware attention mechanism and extend this attention to each feature by considering the heterophily property in graphs.
  • We propose a novel GNN-based recommendation system, called HetroFair, that exploits heterophily to improve fairness.
@@
  • We conduct extensive experiments over six well-known datasets and show that our proposed model outperforms state of the art methods, in terms of both accuracy and fairness metrics. Furthermore, by performing ablation studies, we demonstrate that each component of our proposed method considerably contributes in the performance of the method.
@@ 2.2 Fairness-aware recommendation
The primary objective of popularity debias is to unearth qualified but less interacted items, long-tail ones, which are items that have lower training samples than short-head items. V2HT [17] constructs a heterogeneous item-item graph with four different types of edges to connect long-tails to popular items. SGL [31] brings self-supervised learning to the recommendation setting and creates an auxiliary task for the long-tail recommendation. Zhao et al. [40] disentangle popularity bias into two benign and harmful categories and argue that item quality can also affect popularity bias. They propose a time-aware disentangled framework to identify harmful bias. r-AdjNorm [39] changes the power of the symmetric square root normalization term in graph neural networks to control the normalization process during neighborhood aggregation, aiming for improved results, especially for low-degree items. APDA [42] assigns lower weights to connected edges during the aggregation process and employs residual connections to achieve unbiased and fair representations for users and items in graph collaborative filtering. Chen et al. [7] theoretically analyze the effect of graph convolution on popularity bias amplification and show that stacked multiple graph convolution layers accelerate users’ movement towards popular items in the representation space. Our proposed method takes a different approach to improve items’ side fairness and on the one hand, considers different weights for features during the aggregation process. On the other hand, it directly involves dot products in the normalization process.
@@ 3.3 Light Graph Convolutional
3.4 Heterophily in graphs
Heterophily refers to a property in which nodes of a graph that are not similar to each other, tend to be connected to each other. In other words, it describes the preference for forming connections with nodes that have different characteristics or attributes. This concept is contrasted with homophily, where nodes with similar attributes are more likely to be connected. The homophily rate is a measure used to quantify the tendency of nodes to connect with other nodes that have similar attributes. It can be computed using the following formula [22]:
homophily $\text{homophily rate}=\frac{1}{|V|}\sum_{v\in V}\frac{|\{(w,v):w\in N(v)\wedge C_{v}=C_{w}\}|}{d_{v}},$ (5)
where $C_{x}$ is the class label of node $x$. The range of this metric is within $[0,1]$, where a value of 1 indicates strong homophily, meaning that neighbors between two directions of edges have similar features, in terms of node labels. A value of 00, on the other hand, signifies that neighbor nodes have diverse characteristics [22].
@@ 4.1 Symmetric square root normalization effect
For user $u$, Equation 7 can be written as follows:
@@
In recommendation systems, dot product is a commonly used operator to identify items that are relevant to users. Let $i$ and $i^{\prime}$ be two items that exist in the test profile of user $u$. The prediction scores $y_{ui}$ and $y_{ui^{\prime}}$ are calculated as follows:
$y_{ui}=h_{u}^{\left(k\right)}\cdot h_{i}^{\left(k\right)}=\frac{\sqrt{d_{u}+1}{\sqrt{d_{i}+1}}}{\left(2|E|+|V|\right)^{2}}\left[\sqrt{\left(d_{1}+1\right)}h_{1}^{\left(0\right)}\right.\\ \left.+\ldots+\left(\sqrt{d_{|V|}+1}\right)h_{|V|}^{\left(0\right)}\right]^{2},$ (9)
$y_{ui^{\prime}}=h_{u}^{\left(k\right)}\cdot h_{i^{\prime}}^{\left(k\right)}=\frac{\sqrt{d_{u}+1}{\sqrt{d_{i^{\prime}}+1}}}{\left(2|E|+|V|\right)^{2}}\left[\sqrt{\left(d_{1}+1\right)}h_{1}^{\left(0\right)}\right.\\ \left.+\ldots+\left(\sqrt{d_{|V|}+1}\right)h_{|V|}^{\left(0\right)}\right]^{2}.$ (10)
The only difference between Equations 9 and 10 is the subscripts in the terms $\sqrt{d_{i}+1}$ and $\sqrt{d_{i^{\prime}}+1}$ in the numerator of the Equations. Therefore, if the degree of node $i$ surpasses that of node $i^{\prime}$, it leads to a higher score for node $i$, potentially neglecting the users’ preference for low-degree nodes, despite our knowledge that the user likes both of them. This observation highlights a flaw in the prevalent symmetric square root normalization, employed by graph neural networks, as it tends to exhibit popularity bias and sacrifices items with lower degrees. In an effort to rectify this bias and alleviate its impact on ranking, we propose a novel fairness-aware attention-based model. This model aims to enhance the ranking opportunities for long-tail items, ensuring they have a fair chance of attaining higher positions.
@@ 4.2 Fairness-aware attention
In the previous section, we demonstrated the destructive effect of symmetric square root normalization on the representation of long-tail items. To address this issue, we change light graph convolution, proposed by LightGCN [14], as follows:
@@
where $\sigma\left(\cdot\right)$ is the sigmoid function. Parameter $\lambda$ plays a pivotal role in governing the extent to which the item importance for user $u$ is either amplified or diminished. The sigmoid function transforms the dot product to a normalized range of $\left[0,1\right]$. Therefore, the result of the sigmoid function for long-tail items will be close to zero, and for short-head items, it will be close to one. By incorporating the sigmoid of dot products along with the above-mentioned normalization term, long-tail items achieve a smaller denominator than short-head items. This leads to messages from long-tail items (such as $i$) receiving a higher importance coefficient than those from short-head items (such as $i^{\prime}$). Consequently, this augmentation enhances the similarity between users and long-tail items, compared to the previous case. Conversely, messages from short-head items are propagated with lower coefficients. This not only prevents the over-enlargement of representations, but also ensures that the effects of large representations are gradually transmitted through nodes’ computational graphs.

### v3

@@ Abstract
In recent years, graph neural networks (GNNs) have become a popular tool to improve the accuracy and performance of recommender systems. Modern recommender systems are not only designed to serve end users, but also to benefit other participants, such as items and item providers. These participants may have different or conflicting goals and interests, which raises the need for fairness and popularity bias considerations. GNN-based recommendation methods also face the challenges of unfairness and popularity bias, and their normalization and aggregation processes suffer from these challenges. In this paper, we propose a fair GNN-based recommender system, called HetroFair, to improve item-side fairness. HetroFair uses two separate components to generate fairness-aware embeddings: i) Fairness-aware attention, which incorporates the dot product in the normalization process of GNNs to decrease the effect of nodes’ degrees. ii) Heterophily feature weighting, to assign distinct weights to different features during the aggregation process. To evaluate the effectiveness of HetroFair, we conduct extensive experiments over six real-world datasets. Our experimental results reveal that HetroFair not only alleviates unfairness and popularity bias on the item side but also achieves superior accuracy on the user side. Our implementation is publicly available at https://github.com/NematGH/HetroFair.
@@ 1 Introduction
Graph-based recommendation has become more popular in recent years due to the existence of a bipartite graph in the nature of recommender systems [5, 14, 23, 27, 28, 31]. In this bipartite graph, nodes are divided into two sets: users and items, with edges representing ratings or interactions between them. Graph neural networks (GNNs) are a powerful method for deep learning in graph-structured data, as they can capture the complex and non-Euclidean relationships between the nodes and edges of a graph [43, 9, 11].
With the widespread use of graph neural networks in recommender systems, the competition to build systems with higher accuracy has entered a new phase [14, 8, 18, 36]. Early methods attempted to improve the accuracy of recommender systems to better satisfy end users. However, users are not the only stakeholders, and accuracy is not the sole objective. Recent research demonstrates that recommender systems are multi-stakeholder systems in which, in addition to end users, other participants (especially items and item providers) should also be considered during the recommendation process [1, 2, 4, 20, 21].
@@
On the other hand, the power-law characteristic of the recommendation graphs shows that a large number of items possess a low degree, while only a few of them have a large degree. Figure 1 illustrates the degree distribution of two real-world recommendation datasets. This feature causes a phenomenon known in the literature as popularity bias, where popular items are recommended more frequently than other items. As a result, items with few interactions (long-tail items) are ignored, even if they are of high quality and relevance [3, 19]. Similar investigations have been done for GNN-based recommendation, revealing that long-tail items are often sacrificed during message normalization and aggregation of GNNs [39, 42].
Various methods have been proposed in recent years for long-tail recommendation, which aim to recommend items that belong to the long tail of the distribution of ratings or popularity while simultaneously trying to achieve high accuracy on the user side. Regularization-based approaches attempt to mitigate popularity bias by changing the loss function that penalizes the entire network [44]. However, this often degrades overall performance, as it uniformly restricts the model without considering nuanced variations in item popularity. Another line of research focuses on multi-graph learning, where multiple graphs are optimized for different objectives in an integrated manner [17, 26]. Additionally, causal learning has been leveraged to address popularity bias by constructing causal graphs to analyze the sources of bias and adjust the user-item relevance scores accordingly [29].
Recent studies have shown that message passing in GNNs is a key factor in exacerbating bias against nodes with lower degrees [6, 38]. Some works address this issue by modifying the normalization and aggregation processes in GNN-based recommender systems [39, 42]. However, these approaches overlook two critical aspects: 1) The weights assigned to messages for normalizing the effect of popularity bias are static and non-trainable, remaining fixed across the entire propagation layers. This limits the model’s ability to adaptively adjust for different levels of item popularity. 2) All features (attributes) of a node are treated equally, with a single fixed weight applied to all, despite the fact that different features capture distinct aspects of an item. For example, in movie-related item nodes, some features may represent the number of star actors, others the film producer, and others the genre, each contributing differently to the recommendation process.
@@
In this paper, we propose the HetroFair model to improve item-side fairness in GNN-based recommendation systems. To do so, we first directly address the cause of popularity bias and reduce the effect of nodes’ degrees on enlarging the values of their embeddings by involving dot products during the normalization process. Second, with respect to the heterophily property in graphs, we distinguish between different aspects of an item by considering a specific weight for each feature during the aggregation process. Our key contributions are summarized as follows:
@@
  • We theoretically analyze the effect of message normalization in GNN-based recommendation systems on popularity bias.
@@
  • To generate fair embeddings for users and items, we design a fairness-aware attention mechanism by adding normalization terms derived from the dot product to embeddings.
  • We introduce feature-specific trainable weights to embeddings of users and items, integrated into the message-passing layers of GNNs, to compute heterophily-aware similarity between items and users.
@@
  • We conduct extensive experiments over six well-known datasets and show that our proposed model outperforms state-of-the-art methods in terms of both accuracy and fairness metrics. Furthermore, by performing ablation studies, we demonstrate that each component of our proposed method significantly contributes to the overall performance.
@@ 2.2 Fairness-aware recommendation
The primary objective of popularity debias is to unearth qualified but less interacted items, long-tail ones, which are items that have lower training samples than short-head items. V2HT [17] constructs a heterogeneous item-item graph with four different types of edges to connect long-tails to popular items. SGL [30] brings self-supervised learning to the recommendation setting and creates an auxiliary task for the long-tail recommendation. Zhao et al. [40] disentangle popularity bias into two benign and harmful categories and argue that item quality can also affect popularity bias. They propose a time-aware disentangled framework to identify harmful bias. r-AdjNorm [39] changes the power of the symmetric square root normalization term in graph neural networks to control the normalization process during neighborhood aggregation, aiming for improved results, especially for low-degree items. APDA [42] assigns lower weights to connected edges during the aggregation process and employs residual connections to achieve unbiased and fair representations for users and items in graph collaborative filtering. PopGo [37] proposes a debiasing strategy for collaborative filtering models by estimating interaction-wise popularity shortcuts and adjusting predictions accordingly. It trains a separate shortcut model to learn popularity-related representations and assigns a shortcut degree to each user-item pair. Chen et al. [6] theoretically analyze the effect of graph convolution on popularity bias amplification and show that stacked multiple graph convolution layers accelerate users’ movement towards popular items in the representation space. Our proposed method takes a different approach to improve items’ side fairness and on the one hand, considers different weights for features during the aggregation process. On the other hand, it directly involves dot products in the normalization process.
@@ 3.3 Light Graph Convolutional
We define the size of an embedding $em$ as $\sqrt{em_{1}^{2}+em_{2}^{2}+\cdots+em_{n}^{2}}$, where $n$ is the dimension of $em$ and $em_{i}$ refers to its $i$-th element. This is also known as the norm-2 of $em$. We say embedding $em$ is greater than embedding $em^{\prime}$, denoted by $em>em^{\prime}$, if the size of $em$ is greater than the size of $em^{\prime}$.
@@ 4.1 Symmetric square root normalization effect
Combining Equation 6 with Equation 5 yields the following for user $u$:
@@
In Proposition 1, we exploit this equation to derive a relationship between the embedding sizes of users (and items) and their degrees. Then, in Proposition 2, we derive a relationship between the dot product similarity of the embeddings of items and users and their degrees. Later, in Corollary 1, we combine these two to extract a relationship between the dot product similarities of the embeddings of items and users with their embedding sizes. In recommendation systems, the dot product is a commonly used operator to identify items that are relevant to users.
Proposition 1.
The embedding $h_{u}^{(k)}$ of user $u$ (respectively, the embedding $h_{i}^{(k)}$ of item $i$) is greater than the embedding $h_{u^{\prime}}^{(k)}$ of user $u^{\prime}$ (respectively, the embedding $h_{i^{\prime}}^{(k)}$ of item $i^{\prime}$), if and only if the degree $d_{u}$ of user $u$ (respectively, the degree $d_{i}$ of item $i$) is greater than the degree $d_{u^{\prime}}$ of user $u^{\prime}$ (respectively, the degree $d_{i}^{\prime}$ of item $i^{\prime}$).
Proof.
From Equation 7, the embedding $h_{u}^{(k)}$ for user $u$ is given by:
$h_{u}^{(k)}=\sqrt{d_{u}+1}\times J,$
where $J=\frac{1}{2|E|+|V|}\sum_{v\in V}\sqrt{d_{v}+1}\cdot h_{v}^{(0)}$ is a constant across all users. Similarly, the embedding for user $u^{\prime}$ is:
$h_{u^{\prime}}^{(k)}=\sqrt{d_{u^{\prime}}+1}\times J.$
Since $J$ is constant and $\sqrt{d_{u}+1}$ is a monotonically increasing function of $d_{u}$,
$\sqrt{d_{u}+1}>\sqrt{d_{u^{\prime}}+1}\quad\text{if and only if}\quad d_{u}>d_{u^{\prime}}.$
Thus, $h_{u}^{(k)}>h_{u^{\prime}}^{(k)}$ if and only if $d_{u}>d_{u^{\prime}}$. The same argument applies to item embeddings $h_{i}^{(k)}$ and $h_{i^{\prime}}^{(k)}$, which are proportional to $\sqrt{d_{i}+1}$ and $\sqrt{d_{i^{\prime}}+1}$, respectively.
Proposition 2.
The similarity score between user $u$ and item $i$, computed as the dot product $h_{u}^{(k)}\cdot h_{i}^{(k)}$, is greater than the similarity score between user $u$ and item $i^{\prime}$, represented by $h_{u}^{(k)}\cdot h_{i^{\prime}}^{(k)}$, if and only if the degree $d_{i}$ of item $i$ is greater than the degree $d_{i^{\prime}}$ of item $i^{\prime}$.
Proof.
Inspired by Equations 7, the similarity scores $y_{ui}$ and $y_{ui^{\prime}}$ are given by:
$y_{ui}=h_{u}^{(k)}\cdot h_{i}^{(k)}=\frac{\sqrt{d_{u}+1}\cdot\sqrt{d_{i}+1}}{\left(2|E|+|V|\right)^{2}}\left(\sum_{v\in V}\sqrt{d_{v}+1}\cdot h_{v}^{(0)}\right)^{2},$
$y_{ui^{\prime}}=h_{u}^{(k)}\cdot h_{i^{\prime}}^{(k)}=\frac{\sqrt{d_{u}+1}\cdot\sqrt{d_{i^{\prime}}+1}}{\left(2|E|+|V|\right)^{2}}\left(\sum_{v\in V}\sqrt{d_{v}+1}\cdot h_{v}^{(0)}\right)^{2}.$
Since the term
$\left(\sum_{v\in V}\sqrt{d_{v}+1}\cdot h_{v}^{(0)}\right)^{2}$
is common in both expressions and $\sqrt{d_{u}+1}$ is identical, the comparison of $y_{ui}$ and $y_{ui^{\prime}}$ depends only on the relationship between $\sqrt{d_{i}+1}$ and $\sqrt{d_{i^{\prime}}+1}$. Therefore:
$y_{ui}>y_{ui^{\prime}}\quad\iff\quad\sqrt{d_{i}+1}>\sqrt{d_{i^{\prime}}+1}.$
Since the square root function is monotonically increasing, we have:
$\sqrt{d_{i}+1}>\sqrt{d_{i^{\prime}}+1}\quad\iff\quad d_{i}>d_{i^{\prime}}.$
Thus, $y_{ui}>y_{ui^{\prime}}$ if and only if $d_{i}>d_{i^{\prime}}$. ∎
Corollary 1.
The similarity score between user $u$ and item $i$, represented by the dot product $h_{u}^{(k)}\cdot h_{i}^{(k)}$, is greater than the similarity score between user $u$ and item $i^{\prime}$, $h_{u}^{(k)}\cdot h_{i}^{\prime(k)}$, if and only if the embedding of item $i$ , $h_{i}^{(k)}$, is larger than the embedding of item $i^{\prime}$, $h_{i}^{\prime(k)}$.
Proof.
From Proposition 1, the embedding size $h_{i}^{(k)}$ of an item $i$ increases with its degree $d_{i}$. Proposition 2 states that the similarity score $h_{u}^{(k)}\cdot h_{i}^{(k)}$ is greater than $h_{u}^{(k)}\cdot h_{i^{\prime}}^{(k)}$ if and only if the degree $d_{i}$ of item $i$ is greater than the degree $d_{i^{\prime}}$ of item $i^{\prime}$. Since higher degrees correspond to larger embeddings, it follows that the similarity score $h_{u}^{(k)}\cdot h_{i}^{(k)}$ is greater than $h_{u}^{(k)}\cdot h_{i^{\prime}}^{(k)}$ if and only if the embedding $h_{i}^{(k)}$ is larger than $h_{i^{\prime}}^{(k)}$.
(a) input graph
(b) layer = 1
(c) layer = 2
Figure 2: Illustration of the propagation process in a LightGCN model.
(a) input graph
(b) layer = 1
(c) layer = 2
Figure 3: Illustration of the propagation process in our proposed method.
@@ 4.2 Fairness-aware attention
In the previous section, we demonstrated the destructive effect of symmetric square root normalization on the representation of long-tail items: according to Corollary 1, nodes with larger embeddings have higher degrees and greater similarity scores, which harms fairness and exacerbates popularity bias by disproportionately favoring such nodes. To address this issue and prevent the embedding sizes from growing uncontrollably across multiple layers, it is crucial to control and reduce the embeddings’ size. To achieve this, we modify light graph convolution, proposed by LightGCN [14]. Specifically, since popularity bias arises from high degrees leading to large embeddings, which in turn result in higher similarity scores, this creates a recursive pattern. In our method, we incorporate the similarity scores from the previous layer as a normalization term in the current layer to balance the effects of degree and similarity. The modified version is as follows:
@@
where $\sigma\left(\cdot\right)$ is the sigmoid function. Parameter $\lambda$ plays a pivotal role in governing the extent to which the item importance for user $u$ is either amplified or diminished. The sigmoid function transforms the dot product to a normalized range of $\left[0,1\right]$. Therefore, the result of the sigmoid function for long-tail items will be close to zero, and for short-head items, it will be close to one. By incorporating the sigmoid of dot products along with the above-mentioned normalization term, long-tail items achieve a smaller denominator than short-head items. This ensures that larger embeddings, which are more similar to each other, are divided by larger values and shrink more, while smaller embeddings are divided by smaller values and thus shrink less. Therefore messages from long-tail items (such as $i$) receive a higher importance coefficient than those from short-head items (such as $i^{\prime}$). Consequently, this augmentation enhances the similarity between users and long-tail items, compared to the previous case. Conversely, messages from short-head items are propagated with lower coefficients.
This adjustment leads to changes in the graph structure obtained through successive propagation layers. Figures 2 and 3 clearly illustrate these changes. In Figure 2, related to the LightGCN, we can see that during the application of different propagation layers, popular items establish stronger and closer connections with users, while the connections to less popular items weaken, symbolized by thinner edges. For example, these changes are shown by the thickening of the edges between users and the popular items $i_{4}$, and the thinning of the connections between user $u_{1}$ and item $i_{1}$. As we proceed through layers, this trend intensifies–high-degree items maintain dominance through robust connections, whereas low-degree items experience significantly reduced interaction influence, pushing them further from the users. In contrast, as shown in Figure 3, our proposed method strengthens connections between users and less popular items, making these edges thicker, while connections to popular items become relatively thinner, weakening their influence. Throughout the progression of layers, less popular items retain significant influence with strong connections to users, whereas popular items, though still linked, are pushed farther away from users and are less dominant in the graph structure. This not only prevents the over-enlargement of representations but also ensures that the effects of large representations are gradually transmitted through nodes’ computational graphs.